# Curious Causality-Seeking Agents Learn Meta Causal World

### v2

@@ 5 Related Work
Causal discovery for world understanding provides a rigorous framework for modeling the generative processes that govern complex systems, revealing how one variable brings about changes in another. By explicitly representing cause–effect relationships, causal methods yield more compact and invariant descriptions of reality than do purely statistical correlations (janzing2010causal; li2020causal; yao2022learning; bongers2018causal; NEURIPS2022_cf4356f9; lin2024because). Yet, identifying causal structure from observational data alone is a challenge since multiple directed acyclic graphs can entail the same conditional independencies and v-structures. To resolve these ambiguities, researchers have turned to interventional data—actively perturbing variables to distinguish among candidate graphs (meinshausen2016methods; hyttinen2014constraint; triantafillou2015constraint; brouillard2020differentiable; seitzer2021causal)—and more recent work has even addressed settings with unknown intervention targets (hauser2012characterization; hauser2015jointly; faria2022differentiable; kumar2021disentangling).
Despite these advances, most causal discovery algorithms presume a single uniform, context-independent causal graph (NEURIPS2023_65496a49; alias2021neural; lei2022causal; tomar2021model; zhang2019learning; sontakke2021causal) in the world. Some recent methods have sought to learn context-dependent causal structures (pitis2020counterfactual; chitnis2021camps; wang2023elden; pitis2022mocoda; lowe2022amortized; hwang2023discovery; hwang2024fine; jamshidi2023causal), but they typically require strong prior knowledge (pitis2020counterfactual; huang2020causal) or complete access to a known dynamic model (chitnis2021camps), and they lack active, curiosity-driven exploration strategies for open-ended environments (pearl2009causality).
In contrast, we introduce an intervention-driven framework that empowers an agent to actively probe its surroundings in order to uncover context-dependent causal graphs. By combining targeted interventions with a curiosity-driven exploration policy, our approach adaptively reveals the global causal rules governing rich, previously unseen environments.
@@ 6.1 Experiment Setup
We compare our method against several baselines: MLP, a dense model that predicts transition dynamics $p(x_{t+1}|x_{t},a_{t})$; Modular, a modular network with separate modules for each state variable; GNN (kipf2019contrastive), a graph neural network for predicting transition dynamics; NPS (alias2021neural), which learns sparse and modular dynamics; CDL (wang2022causal), which learns a static causal model from data; GRADER (ding2022generalizing), which learns a static causal model using conditional independence tests; NCD (hwang2024fine), a neural causal discovery algorithm for learning causal graphs for each sample; and FCDL (hwang2024fine) learns causal graph to facilitate the robustness for reinforcement learning. Please check Appendix B for details of the experimental setup.
@@ 6.3 Experiment Results
Downstream Task We evaluate our method against baselines on the Chemical downstream task, using the trained model to predict transition dynamics for model-based reinforcement learning. As shown in Table 2, our approach achieves the highest reward, demonstrating its effectiveness in learning a more accurate causal world model.
@@ 7 Conclusion, Limitation and Future Work
Limitation and Future Work Despite these advances, real-world exploration remains subject to inherent constraints: certain states are fundamentally unreachable via any sequence of interventions, and the implications of such unobservability for causal inference have yet to be analyzed (see Appendix A.4.1). Moreover, interventions incur costs, and practical agents must operate within limited budgets. Future research will investigate the impact of unreachable states on the identifiability and accuracy of learned causal structures, and will develop budget-aware exploration strategies to optimally allocate limited intervention resources.
@@ A.2 Swap-Label Equivalence
(a) Meta-Causal Graph with two meta states $u_{1}$ and $u_{2}$ and corresponding causal subgraph $M_{u_{1}}$ and $M_{u_{2}}$.
(b) Meta-Causal Graph with two meta states $\hat{u}_{1}$ and $\hat{u}_{2}$ and corresponding causal subgraph $\hat{M}_{\hat{u}_{1}}$ and $\hat{M}_{\hat{u}_{2}}$.
@@ A.3 Proof of Theorem 1
[…]
@@ A.4 Observational Equivalence
(a) Meta-Causal Graph with two meta states.
(b) Meta-Causal Graph with three meta states.
@@ A.5 Proof of Theorem 2
(a) Ground truth graph.
(b) Causal skeleton matrix of the ground truth graph.
(c) Learned graph with overparameterization.
(d) Causal skeleton matrix of the learned graph.
Figure 8: Illustration of the Theorem 2. The upper figure shows the ground truth graph and its causal skeleton matrix. The lower figure shows the learned graph with overparameterization and its causal skeleton matrix.
@@ A.6 Disscussion on Assumption 1
[…]
@@
(a) Causal graph with 3 meta states
(b) Learned graph with 3 meta states
(c) Causal graph with 2 meta states
(d) Learned graph with 2 meta states
Figure 9: Illustration of the Assumption 1. The upper figure shows the ground truth causal graph with three meta states and the corresponding causal skeleton matrix. The lower figure shows the learned graph with two meta states and the corresponding causal skeleton matrix.
@@ A.8 Reachability of Interventions
A state corresponding to index $i$ is reachable from initial state $\mathbf{z}_{0}$ if and only if there exists $k\geq 0$ such that $[(TF)^{k}\mathbf{z}_{0}][i]>0$.
@@
($\Rightarrow$): If state $i$ is reachable from $\mathbf{z}_{0}$, then by definition there exists a finite sequence of intervention-transition steps that leads from $\mathbf{z}_{0}$ to a state where the $i$-th component is active. This sequence corresponds to some power $k$ of the operator $TF$, hence $[(TF)^{k}\mathbf{z}_{0}][i]>0$.
($\Leftarrow$): If $[(TF)^{k}\mathbf{z}_{0}][i]>0$ for some $k\geq 0$, then by the definition of matrix-vector multiplication, there exists a computational path through $k$ iterations of intervention-transition operations that activates the $i$-th state component, demonstrating reachability. ∎
@@
The state corresponding to index $i$ is feasible from initial state $\mathbf{z}_{0}$ by intervention if and only if there exists $k\geq 0$ such that $[F(TF)^{k}\mathbf{z}_{0}][i]>0$.
@@
Reachability Analysis: Starting from $\mathbf{z}_{0}=[1,0,0,0]^{\top}$ (state (0,0)(0,0)( 0 , 0 )):
@@ B.1 Environment
These settings test our agent’s ability to: (1) discover the latent meta states (root node colors) that determine causal structure transitions, (2) learn the corresponding causal subgraphs through interventional exploration, and (3) generalize to unseen state configurations during evaluation. During the test, some nodes are corrupted with noise. The agent needs to learn the causal graph to match the colors of nodes to the target.
Magnetic
@@ B.3 Environment Configurations
B.4 Extended Results
@@
B.5 Compute Resources and Environment Details
@@
B.6 Code and Demo
The code is available at https://anonymous.4open.science/r/MCG-653D, and demonstrations can be found at https://sites.google.com/view/meta-causal-world.
@@ C.1 World Models
Within the first trajectory, numerous approaches leverage world models as predictive mechanisms to facilitate planning processes. Notable exemplars include methodologies employing Monte Carlo Tree Search (MCTS) to identify optimal action sequences [schrittwieser2020mastering, dedieu2025improving] and techniques utilizing cross-entropy methods to efficiently sample continuous actions [hansen2023td]. Complementary to these, a substantial corpus of research utilizes world models to augment policy learning, wherein the learned dynamics models either provide supplementary supervision signals or function as synthetic environments for policy optimization. This subcategory encompasses seminal works such as Dreamer [hafner2019dream, hafner2020mastering, hafner2023mastering], SimPLe [kaiser2019model], IRIS [micheli2022transformers], $\Delta$-IRIS [micheli2024efficient], and DART [agarwal2024learning]. These methodologies predominantly employ model-based reinforcement learning frameworks, wherein learned dynamics models generate synthetic data to facilitate policy model training.

### v3

@@ 5 Related Work
Causal Discovery for Open-Ended World.
Causal discovery provides a rigorous framework for modeling the generative processes that govern complex systems, revealing how one variable brings about changes in another. By explicitly representing cause–effect relationships, causal methods yield more compact and invariant descriptions of reality than do purely statistical correlations (janzing2010causal; li2020causal; yao2022learning; bongers2018causal; chen2024doubly; cheng2025empowering; NEURIPS2022_cf4356f9; lin2024because). Causal representation learning provides a useful tool to learn causal structure, which aims to recover causally meaningful latent factors and structures, enabling invariance and transfer in downstream tasks (scholkopf2021toward; locatello2020weakly; ke2021systematic; Yang_2021_CVPR; NEURIPS2023_fc657b7f; 10.1145/3459637.3482305). However, most of the work of causal representation learning focusses on learning a static causal structures passively from observational data.
However, identifying causal structure from observational data alone is a challenge since multiple directed acyclic graphs can entail the same conditional independencies and v-structures (wang2024building). To resolve these ambiguities, researchers have turned to the collection of active causal interventional data by actively perturbing variables to distinguish among candidate graphs (meinshausen2016methods; hyttinen2014constraint; triantafillou2015constraint; brouillard2020differentiable; seitzer2021causal). More recent work has even addressed settings with unknown intervention targets (hauser2012characterization; hauser2015jointly; faria2022differentiable; kumar2021disentangling).
Despite these advances, most causal discovery algorithms presume a single uniform, context-independent causal graph (NEURIPS2023_65496a49; alias2021neural; lei2022causal; tomar2021model; zhang2019learning; sontakke2021causal; yu2024learning; cao2025towards) in the world. Some recent methods have sought to learn causal structures in changing world (pitis2020counterfactual; chitnis2021camps; wang2023elden; pitis2022mocoda; lowe2022amortized; hwang2023discovery; hwang2024fine; jamshidi2023causal), but they typically require strong prior knowledge (pitis2020counterfactual; huang2020causal) or complete access to a known dynamic model (chitnis2021camps), and they lack active, curiosity-driven exploration strategies for open-ended environments (pearl2009causality).
We introduce an intervention-driven framework that empowers an agent to actively probe its surroundings in order to uncover context-dependent causal graphs. By combining targeted interventions with a curiosity-driven exploration policy, our approach adaptively reveals the global causal rules governing rich, previously unseen environments. Extended related work could be found in Appendix C.
@@ 6.1 Experiment Setup
We compare our method against several baselines: MLP, a dense model that predicts transition dynamics $p(x_{t+1}|x_{t},a_{t})$; Modular, a modular network with separate modules for each state variable; GNN (kipf2019contrastive), a graph neural network for predicting transition dynamics; NPS (alias2021neural), which learns sparse and modular dynamics; CDL (wang2022causal), which learns a static causal model from data; GRADER (ding2022generalizing), which learns a static causal model using conditional independence tests; Sandy-Mixure (pitis2020counterfactual), which uses Jacobian matrices of MLP layers to learn the local causal graph; Transformer (urpi2024causal), which infers causal graphs by attention; NCD (hwang2024fine), a neural causal discovery algorithm for learning causal graphs for each sample; and FCDL (hwang2024fine) learns causal graph to facilitate the robustness for reinforcement learning. Please check Appendix B for details of the experimental setup.
@@ 6.3 Experiment Results
Downstream Task We evaluate our method against baselines on the Chemical downstream task, using the trained model to predict future states with the learned model for decision making. As shown in Table 2, our approach achieves the highest reward, demonstrating its effectiveness in learning a more accurate causal world model.
@@
Partial Observation We modify the Chemical environment to explicitly simulate challenges arising from limited observation windows. This experimental setup evaluates how our method performs as new variables gradually become observable. The results indicate that our approach remains robust under partial observability and shows strong potential for handling more realistic, dynamically evolving environments. The results are shown in Table 8.
Efficiency We compare baselines with longer training rounds and more parameters to show efficiency of our method. Results are reported in Table 10 and Table 11.
@@ 7 Conclusion, Limitation and Future Work
Limitation and Future Work Despite these advances, real-world exploration remains subject to inherent constraints: certain states are fundamentally unreachable via any sequence of interventions, and the implications of such unobservability for causal inference have yet to be analyzed (see Appendix A.8). Moreover, interventions incur costs, and practical agents must operate within limited budgets. Future research will investigate the impact of unreachable states on the identifiability and accuracy of learned causal structures, and will develop budget-aware exploration strategies to optimally allocate limited intervention resources.
@@ A.2 Swap-Label Equivalence
(a) Two meta states $(u_{1},u_{2})$ with corresponding subgraphs $M_{u_{1}}$ and $M_{u_{2}}$.
(b) Label-swapped meta states $(\hat{u}_{1},\hat{u}_{2})$ with corresponding subgraphs $\hat{M}_{\hat{u}_{1}}$ and $\hat{M}_{\hat{u}_{2}}$.
@@ A.3 Proof of Theorem 1
Moreover, since $M_{u_{1}}\neq M_{u_{2}}$, there exists at least one position $(i,j)$ where $M_{u_{1}}[i,j]\neq M_{u_{2}}[i,j]$.
@@ A.4 Observational Equivalence
(a) Two meta states: partition (left) and Meta-Causal Graph (right).
(b) Three meta states: partition (left) and Meta-Causal Graph (right).
@@ A.5 Proof of Theorem 2
(a) Ground truth: graph (left) and skeleton matrix (right).
(b) Over-parameterized: learned graph (left) and skeleton matrix (right).
Figure 8: Illustration of Theorem 2. Each pair places the graph (left) alongside its causal skeleton matrix (right): (a) ground truth; (b) overparameterized solution.
@@ A.6 Disscussion on Assumption 1
where $\mathbb{I}[\cdot]$ is the indicator function, ensuring that an edge $(i,j)$ exists in the learned graph if and only if it exists in at least one contributing subgraph.
@@
(a) Three meta states: causal graph (left) and learned graph (right).
(b) Two meta states: causal graph (left) and learned graph (right).
Figure 9: Illustration of the Assumption 1. The left figure shows the ground truth causal graph with three meta states and the corresponding causal skeleton matrix. The right figure shows the learned graph with two meta states and the corresponding causal skeleton matrix.
@@ A.7 Proof of Proposition 1
Remark 1.
Score-based causal discovery methods identify structures only up to a Markov equivalence class, since different directed graphs can encode the same set of conditional independencies. Interventions, however, can break this equivalence: by actively intervening on variables, one can distinguish among graphs within the same class.
@@ A.8 Reachability of Interventions
A state corresponding to index $i$ is reachable from initial state $\mathbf{z}_{0}$ if and only if there exists $k\geq 0$ such that $(TF)^{k}\mathbf{z}_{0}][i]>0$.
@@
($\Rightarrow$): If state $i$ is reachable from $\mathbf{z}_{0}$, then by definition there exists a finite sequence of intervention-transition steps that leads from $\mathbf{z}_{0}$ to a state where the $i$-th component is active. This sequence corresponds to some power $k$ of the operator $TF$, hence $(TF)^{k}\mathbf{z}_{0}][i]>0$.
($\Leftarrow$): If $(TF)^{k}\mathbf{z}_{0}][i]>0$ for some $k\geq 0$, then by the definition of matrix-vector multiplication, there exists a computational path through $k$ iterations of intervention-transition operations that activates the $i$-th state component, demonstrating reachability. ∎
@@
The state corresponding to index $i$ is feasible from initial state $\mathbf{z}_{0}$ by intervention if and only if there exists $k\geq 0$ such that $F(TF)^{k}\mathbf{z}_{0}][i]>0$.
@@
Reachability Analysis: Starting from $\mathbf{z}_{0}=[1,0,0,0]^{\top}$ (state $(0,0)$):
@@
A.9 Action as Intervention
Our method identifies context-dependent causal graphs (meta-graphs) via interventions. Actions are how the curiosity-driven agent realizes those interventions. When actions can directly intervene each state (e.g., a chemical environment), we deploy targeted interventions to isolate edges and quickly distinguish meta states and their subgraphs. When not all interventions are feasible, we rely on the reachability assumption.
Verification Results.
We add a new experiment based on a manipulation task to demonstrate that our method can handle cases where actions cannot intervene on all variables. We also modify this environment to test generalization. The results show that our method maintains strong performance and generalizes well, even when only a subset of variables is directly intervenable.
Table 4: Performance on partial-intervention manipulation tasks.
Reward Small Magnetic Force ($\times 0.02$) High Ball Density ($\times 10$) Extra Table Friction
MLP 4.011 $\pm$ 0.030 3.999 $\pm$ 0.038 3.886 $\pm$ 0.241
FCDL 4.482 $\pm$ 0.622 4.451 $\pm$ 0.657 4.193 $\pm$ 0.655
MCG 6.173 $\pm$ 0.255 5.083 $\pm$ 0.245 5.517 $\pm$ 0.134
A.10 Lossy Representation and Misclassification Probability
We analyze how lossy representation affects the probability of mapping samples to prototype embeddings. Table 5 summarizes the key notations used in this derivation.
Table 5: Notation summary.
Symbol Meaning
$U=u$ Ground-truth meta states
$\hat{U}=\hat{u}_{k}$ Prototype (codebook) embeddings
$C(x)$ True meta state of sample $x$
$\hat{C}(x)$ Codebook entry selected by encoder
$\mu_{u}$ Prior probability of state $x$ with true meta state $u$
$p_{k}=\sum_{u\in U}\mu_{u}p_{k}^{u}$ Probability of mapping sample $x$ to codebook entry $e_{\hat{u}_{k}}$
We assume that for each state $x$ with true meta state $C(x)=u$, the probability that it is mapped by the encoder to the codebook entry index $\hat{u}_{k}$ is $p^{u}_{k}=P(\hat{C}(x)=\hat{u}_{k}\mid C(x)=u)$. This probability reflects the representation power and discriminability of the encoder, indicating how likely samples of the same true meta state are mapped to different codebook entry.
From Definition 5, $\hat{C}(x)$ and $C(x)$ are observationally equivalent if, for all $\hat{u}_{k}\in\hat{U}$ and all $x_{1},x_{2}\in\{x\mid\hat{C}(x)=\hat{u}_{k}\}$, it holds that $C(x_{1})=C(x_{2})$.
The probability that $\hat{C}(x)$ and $C(x)$ are observationally equivalent is given by
$P(C(x)=u,\forall x^{\prime}\in x\mid\hat{C}(x)=\hat{u}_{k},C(x^{\prime})=u)=(1-p_{k}+p^{u}_{k}\mu_{u})^{n},$
where $n$ is the number of samples.
Hence, the overall probability of observational equivalence is
$P_{\text{equiv}}=\sum_{k}p_{k}\sum_{u\in U}\mu_{u}(1-p_{k}+p^{u}_{k}\mu_{u})^{n}.$
The probability that all samples assigned to the same codebook entry index $\hat{C}(x)$ belong to different true meta states is
$P_{\text{diff}}=\sum_{k}\sum_{u\in U}(p_{k}-\mu_{u}p^{u}_{k})(1-p_{k}+p^{u}_{k}\mu_{u})^{n}.$
The probability of misclassification, i.e., that a codebook entry corresponds to mixed true meta states, is
$P_{\text{misclass}}=1-\sum_{k}p_{k}\sum_{u\in U}(1-p_{k}+p^{u}_{k}\mu_{u})^{n}.$
Considering lossy representation, the misclassification probability simplifies to
$P(\text{misclassification})=1-\sum_{k}\sum_{u\in U}\mu_{u}p^{u}_{k}(1-p_{k}+p^{u}_{k}\mu_{u})^{n}.$
If $n(p_{k}+p^{u}_{k}\mu_{u})\ll 1$, we approximate
$P(\text{misclassification})\approx n\sum_{u}\sum_{k}\mu_{u}p^{u}_{k}(p_{k}-p^{u}_{k}\mu_{u})=\sum_{k}\left[p_{k}^{2}-\sum_{u}(\mu_{u}p^{u}_{k})^{2}\right].$
In a more accurate representation, for each codebook entry index $k$, the distribution of $\mu_{u}p^{u}_{k}$ concentrates on a single true state—one $p^{u}_{k}$ increases while others decrease, thereby increasing $\sum_{u}(\mu_{u}p^{u}_{k})^{2}$ and reducing each term $\left[p_{k}^{2}-\sum_{u}(\mu_{u}p^{u}_{k})^{2}\right]$. Consequently, higher representation accuracy monotonically decreases the overall misclassification probability, reaching zero when each prototype corresponds purely to one true meta state.
@@ B.1 Environment
These settings test our agent’s ability to: (1) discover the latent meta states (root node colors) that determine causal structure transitions, (2) learn the corresponding causal subgraphs through interventional exploration, and (3) generalize to unseen state configurations during evaluation. During the test, some nodes are corrupted with noise. The agent needs to learn the causal graph to match the colors of nodes to the target. The agent starts from a random color configuration and must transform a 10-node graph to match a target color pattern. Actions intervene on nodes, changing their color and that of all their descendants according to the hidden causal graph. The episode reward is the negative Hamming distance to the target (0 for a perfect match).
Magnetic.
@@ B.3 Environment Configurations
B.4 Partial Observability
Table 8: Prediction accuracy under partial observability in the Chemical environment.
Prediction Accuracy $n=2$ $n=4$ $n=6$
MLP 31.93 $\pm$ 0.16 32.47 $\pm$ 1.61 33.72 $\pm$ 2.11
FCDL 63.75 $\pm$ 13.75 53.89 $\pm$ 11.95 52.11 $\pm$ 16.43
MCG (ours) 73.52 $\pm$ 9.56 63.71 $\pm$ 5.79 57.05 $\pm$ 14.63
B.5 Extended Results
@@
B.6 Compute Resources and Environment Details
@@
B.7 Code and Demo
B.8 Prediction accuracy for three intrinsic terms
Table 9: Prediction accuracy for three intrinsic terms. All values are reported in percentage $\%$.
Intrinsic Term Fork Chain
n=2n{=}2 n=4n{=}4 n=6n{=}6 n=2n{=}2 n=4n{=}4 n=6n{=}6
Prediction Uncertainty $52.25\pm 16.73$ $59.46\pm 22.38$ $52.25\pm 28.46$ $59.52\pm 8.82$ $54.81\pm 5.47$ $59.28\pm 5.03$
Feature Discrepancy $46.11\pm 8.38$ $41.17\pm 5.67$ $43.67\pm 2.47$ $56.01\pm 13.13$ $51.88\pm 7.22$ $58.27\pm 8.53$
Predictive-Distribution Discrepancy $54.23\pm 8.72$ $46.27\pm 9.74$ $43.92\pm 1.85$ $36.83\pm 3.11$ $40.33\pm 4.89$ $45.40\pm 2.83$
[…]
B.9 Efficiency
We compare other methods (GNN and MLP) with longer training rounds and more parameters to show efficiency of our method. Results are reported in Table 10 and Table 11.
Table 10: Comparison under the same model size. We evaluate the efficiency of our method by comparing MCG with GNN and MLP trained for longer rounds and with larger parameter counts.
Acc n=2 Acc n=4 Acc n=6 Reward n=2 Reward n=4 Reward n=6
GNN (30k round) 38.13 25.81 20.75 6.83 6.06 6.83
GNN (25k round) 36.37 26.55 20.73 6.84 5.98 6.19
MLP (30k round) 30.47 28.94 27.84 6.20 6.13 6.71
MLP (25k round) 31.47 30.32 30.42 6.18 6.03 6.23
MCG (10k round) 47.01 48.37 46.35 11.96 10.83 10.70
MCG (12k round) 48.30 48.40 47.28 12.45 10.77 10.84
MCG (15k round) 71.56 59.17 49.58 14.65 14.06 13.28
Table 11: Comparison under the same number of training rounds. MCG is compared with GNN and MLP models trained for the same number of rounds but with different model sizes (parameter counts shown in parentheses).
Acc n=2 Acc n=4 Acc n=6 Reward n=2 Reward n=4 Reward n=6
MLP (parameter $\times 3.71$) 24.65 25.05 23.71 6.29 6.50 7.93
GNN (parameter $\times 1.66$) 37.58 30.61 30.05 6.34 6.49 7.90
MCG (parameter $\times 1$) 71.56 59.17 49.58 14.65 14.06 13.28
@@ C.1 World Models
Within the first trajectory, numerous approaches leverage world models as predictive mechanisms to facilitate planning processes. Notable exemplars include methodologies employing Monte Carlo Tree Search (MCTS) to identify optimal action sequences [schrittwieser2020mastering, dedieu2025improving, feng2023chessgpt] and techniques utilizing cross-entropy methods to efficiently sample continuous actions [hansen2023td]. Complementary to these, a substantial corpus of research utilizes world models to augment policy learning, wherein the learned dynamics models either provide supplementary supervision signals or function as synthetic environments for policy optimization. This subcategory encompasses seminal works such as Dreamer [hafner2019dream, hafner2020mastering, hafner2023mastering], SimPLe [kaiser2019model], IRIS [micheli2022transformers], $\Delta$-IRIS [micheli2024efficient], and DART [agarwal2024learning]. These methodologies predominantly employ model-based reinforcement learning frameworks, wherein learned dynamics models generate synthetic data to facilitate policy model training.